# Jarne Verhaeghe IDLab

### v1

@@ Acknowledgments
This research was funded by the FWO Junior Research project HEROI2C which investigates hybrid machine learning for improved infection management in critically ill patients (Ref. G085920N). Jarne Verhaeghe is funded by the Research Foundation Flanders (FWO, Ref. 1S59522N). Part of this research was supported through the Flemish Government (AI Research Program).
Appendix A Synthetic Data
A.1 Setup 1
For setup 1, inspired by Wu et al. (2024), six independent covariates are sampled from various distributions representing both continuous and discrete values:
$X_{1},X_{2},X_{3},X_{4}$ $\sim\text{Normal}(0,1)$
[…]
$X_{6}$ $\sim\text{Uniform}(-3,3)$
Scenario Treatment function
$3$ $T_{\mu}=-0.8+X_{1}+0.1X_{2}-0.1X_{3}+0.2X_{4}+0.1X_{5}+0.1X_{6}+\frac{3}{2}X_{3}^{2}$
Table 2: The treatment functions for all scenarios in setup 1.
Scenario Treatment $T$ Treatment Assignment Distribution
1 $179T_{\mu}+$ 17 $\text{Normal}(0,5)$
2 $2215T_{\mu}+$ 22 $\text{StudentT}(df=2)$
3 $159T_{\mu}+$ 15 $\text{Normal}(0,5)$
4 $649\frac{e^{T_{\mu}}}{1+e^{T_{\mu}}}-$ 6 $\text{Normal}(0,5)$
5 $1842\frac{1}{1+e^{T_{\mu}}}+$ 18 $\text{Normal}(0,5)$
6 $137log(|T_{\mu}|+0.001)+$ 13 $\text{Normal}(0,4)$
7 $167T_{\mu}+$ 16 $\text{Normal}(0,1)$
8 $167T_{\mu}+$ 16 $20\cdot\text{Beta}(\alpha=2,\beta=8)$
Table 3: The propensity functions per scenario for Setup 1
$Y$ $\sim-1-(2X_{1}+2X_{2}+3X_{3}^{3}-20X_{4}-2X_{5}+20X_{6})$
A.2 Setup 2
Setup 2 tests the different treatment assignment distributions in the two different scenarios, which is the same experimental setup as proposed by Schröder et al. (2024). The covariates are sampled from a discrete uniform distribution. The treatment is sampled from the treatment assignment distributions shown in Table 4. The outcome function is sampled from a normal distribution with a mean determined by a sinus function based on both $X$ and $T$:
[…]
$Y$ $\sim sin\left((0.05\pi)(T-X)\right)+\text{Normal}(0,0.1)$
Scenario Treatment Assignment Distribution
2 $T\sim\text{Normal}(5X,10)$
Table 4: The propensity functions per scenario for Setup 2
Appendix B Propensity Distribution Estimation
@@
Algorithm 1 presents the propensity distribution estimation using Conformal Predictive Systems (CPS). This results in a propensity distribution array $\pi_{arr}$ with the calculated propensity density for each sample in $X_{cal}$. $exp$ is the exponential function and $len(X)$ denotes the length of the array $X$.
@@
1:  Input: training covariates $X_{tr}$, calibration covariates $X_{cal}$, training treatment values $T_{tr}$, calibration treatment values $T_{cal}$, Kernel Density Estimator $KD$
@@
3:  calibrate CPS on $X_{cal}$ and $T_{cal}$
4:  Define $\pi_{arr}$ with length $len(X_{cal})$
@@
9:  return $\pi_{arr}$
@@
In comparison to the work of Schröder et al. (2024), our approach differs in several key aspects. First, the aim of their work is different from ours. The aim of Schröder et al. (2024) is to provide prediction intervals for the causal effect of treatment interventions where the treatment value is continuous. In our work, the goal is to provide prediction intervals for dose-response models instead of treatment interventions, answering a different causal question. However, adjusting our work to interventions is possible; In the case of soft interventions, the target distribution propensity changes and thus substituting the current uniform distribution in the weights $w(x)$ with the new target propensity distribution covers the soft intervention case. For hard interventions, this is an evaluation for a single treatment value which is similar to the local propensity method, but for only that target treatment value. Secondly, their approach differs in their conformal prediction approach where they want to provide correct prediction intervals for a single sample, single $\alpha$ value, and single treatment using a mathematical solver based on the proposed weighted conformal prediction by Gibbs & Candes (2021). Thirdly, they frame the propensity or covariate shift differently as either a Dirac distribution for a hard intervention, or a different propensity distribution in the case of a soft intervention. This is a direct consequence of their aim to quantify the causal effect of a single intervention, compared to providing a dose-response model in our case which requires a uniform assumption. Fourthly, the experimental setup of Schröder et al. (2024) does not address the impact of a treatment covariate shift as shown by Figure 4 and Figure 5 where even standard conformal prediction (CP) achieves the required empirical coverage. Lastly, we also approach the propensity estimation in cases with unknown propensity as an uncertainty quantification problem and tackle it with conformal predictive systems. In the end, our approach offers a different solution on continuous treatment effects through dose-response modelling.

### v2

@@ 8 Conclusion
This research was funded by the FWO Junior Research project HEROI2C which investigates hybrid machine learning for improved infection management in critically ill patients (Ref. G085920N). Jarne Verhaeghe is funded by the Research Foundation Flanders (FWO, Ref. 1S59522N). Jef Jonkers is funded by the Research Foundation Flanders (FWO, Ref. 1S11525N). Part of this research was supported through the Flemish Government (AI Research Program).
Data and code availability
All code, data, and additional results are available open-source on https://github.com/predict-idlab/dose-response-conformal-prediction.
Author Contribution
Jarne Verhaeghe conceived the study design, wrote and revised the manuscript, and performed the experiments. Jef Jonkers helped conceive the study, and helped write and review the manuscript, Sofie Van Houcke supervised the study and helped review the manuscript.
Declaration of generative AI and AI-assisted technologies in the writing process
During the preparation of this work, the author(s) used ChatGPT and Gemini to optimize text and improve readability. After using this tool/service, the author(s) reviewed and edited the content as needed and take(s) full responsibility for the content of the published article.
Appendix A Finite Sample Coverage Guarantees
For counterfactual prediction intervals, the ideal goal is to achieve the following general conditional coverage guarantee:
$\mathbb{P}_{Y\sim P_{Y|T=t,X=x}}(Y(t)\in\hat{C}(x,t)|X=x)\geq 1-\alpha,\text{where }t\in[t_{L},t_{U}]$ (29)
which, under the strong ignorability assumption, is equivalent to:
$\mathbb{P}_{Y\sim P_{Y|T=t,X=x}}(Y\in\hat{C}(x,t)|X=x,T=t)\geq 1-\alpha.$ (30)
However, constructing non-trivial prediction intervals with such conditional guarantees is generally impossible without additional modelling assumptions, as shown in foygel_barber_limits_2021. Even under the relaxed conditional guarantee, where conditioning is only on the treatment value, as in binary treatment settings (lei_conformal_2021):
$\mathbb{P}_{Y\sim P_{X}\times P_{Y|T=t,X}}(Y\in\hat{C}(X,t)|T=t)\geq 1-\alpha,$ (31)
the problem persists when the treatment variable $t$ is continuous.
A.1 Proposed Framework
To address this challenge, we introduce a distribution shift in the treatment variable by moving from the generalized propensity distribution to a user-specified interventional distribution, $T_{n+1}\sim\tilde{P}_{T}$. We then leverage the weighted conformal prediction (WCP) framework to construct prediction intervals. This approach allows us to build on prior theoretical coverage results under both oracle and estimated likelihood functions (tibshirani_conformal_2019; lei_conformal_2021).
Table 5 outlines the two interventional distributions utilized in this work: global propensity, local propensity, and $\delta$-propensity (Dirac delta). The latter corresponds to a hard intervention. Relaxing the $\delta$-propensity to the local propensity enables the construction of non-trivial prediction intervals (see Remark 4). Notably, when $T\in\{0,1\}$, our approach under $\delta$-propensity aligns with the counterfactual inference framework for binary treatments proposed in lei_conformal_2021. Table 6 shows the translation to a uniform interventional distribution where every dose within range is equally likely to be intervened on.
General Global propensity Local propensity $\delta$-propensity
$\tilde{P}_{T|X}$ $\tilde{P}_{T}$ $\frac{f_{\tilde{P}_{T}}(T,X)K(\frac{T-t}{h})}{\int_{t_{L}}^{t_{U}}f_{\tilde{P}_{T}}(T,X)K(\frac{T-t}{h})dT}$ $\delta(T-t)$
$w(X,T)$ $\frac{f_{\tilde{P}_{T}}(T,X)}{\pi(T|X)}$ $\frac{f_{\tilde{P}_{T}}(T,X)K\left(\frac{T-t}{h}\right)}{\pi(T|X)}$ $\frac{\delta(T-t)}{\pi(T|X)}$
$\hat{w}(X,T)$ $\frac{f_{\tilde{P}_{T}}(T,X)}{\hat{\pi}(T|X)}$ $\frac{f_{\tilde{P}_{T}}(T,X)K\left(\frac{T-t}{h}\right)}{\hat{\pi}(T|X)}$ $\frac{\delta(T-t)}{\hat{\pi}(T|X)}$
Table 5: Translation of general interventional distribution framework to WCP global, local, and $\delta$-propensity.
General Global propensity Local propensity
$\tilde{P}_{T|X}$ $Uniform(t_{L},t_{U})$ $\frac{\mathbbm{1}_{[t_{L},t_{U}]}(T)K(\frac{T-t}{h})}{\int_{t_{L}}^{t_{U}}\mathbbm{1}_{[t_{L},t_{U}]}(T)K(\frac{T-t}{h})dT}$
$w(X,T)$ $\frac{\mathbbm{1}_{[t_{L},t_{U}]}(T)}{\pi(T|X)}$ $\frac{\mathbbm{1}_{[t_{L},t_{U}]}(T)K\left(\frac{T-t}{h}\right)}{\pi(T|X)}$
$\hat{w}(X,T)$ $\frac{\mathbbm{1}_{[t_{L},t_{U}]}(T)}{\hat{\pi}(T|X)}$ $\frac{\mathbbm{1}_{[t_{L},t_{U}]}(T)K\left(\frac{T-t}{h}\right)}{\hat{\pi}(T|X)}$
Table 6: Example of a uniform interventional distribution for WCP global and local.
A.2 Proposition: Finite-Sample Guarantees
Proposition 1. following tibshirani_conformal_2019; lei_conformal_2021 Assume $(X_{i},T_{i},Y_{i})\overset{i.i.d.}{\sim}P_{X}\times P_{T|X}\times P_{Y|T,X}$, i=1,,ni=1,...,n; the likelihood ratio $w(X,T)\propto\frac{d\tilde{P}_{T|X}}{dP_{T|X}}$; and the estimated likelihood ratio $\hat{w}(X,T)$. Using WCP to construct $\hat{C}(X,T)$, the following finite-sample bounds apply:
  • S1.
    (Oracle Likelihood Ratio) If $\hat{w}(\cdot,\cdot)=w(\cdot,\cdot)$, i.e. oracle likelihood ratio function; then,
    $1-\alpha\leq\mathbb{P}_{(X,T,Y)\sim P_{X}\times\tilde{P}_{T|X}\times P_{Y|T,X}}\{Y\in\hat{C}(X,T)\}$ (32)
  • S2.
    (Finite Sample with Regularity Conditions) If $\hat{w}(\cdot,\cdot)=w(\cdot,\cdot)$; the non-conformity scores $S_{i}$ have no ties almost surely; $\tilde{P}_{T|X}\times P_{X}$ is absolutely continuous with respect to $P_{T|X}\times P_{X}$; and $(\mathbb{E}_{(X,T)\sim P_{X}\times P_{T|X}}[w(X,T)^{r}])^{\frac{1}{r}}\leq M_{r}<\infty$ where $r>0$ and $M_{r}$ denotes the upper bound of the $r$-th moment of the likelihood ratio; then,
    $1-\alpha\leq\mathbb{P}_{(X,T,Y)\sim P_{X}\times\tilde{P}_{T|X}\times P_{Y|T,X}}\{Y\in\hat{C}(X,T)\}\leq 1-\alpha+cn^{\frac{1}{r-1}}$ (33)
    where $c$ is an arbitrary positive constant depending on $M_{r}$ and $r$.
  • S3.
    (Estimated Likelihood Ratio) If $\hat{w}(\cdot,\cdot)\neq w(\cdot,\cdot)$; $\Delta_{w}=\frac{1}{2}\mathbb{E}_{(X,T)\sim P_{X}\times P_{T|X}}[|\hat{w}(X,T)-w(X,T)|]$; $(\mathbb{E}_{(X,T)\sim P_{X}\times P_{T|X}}[\hat{w}(X,T)^{r}])^{\frac{1}{r}}\leq M_{r}<\infty$; and further assuming the same assumptions as in S2.; then,
    $1-\alpha-\Delta_{w}\leq\mathbb{P}_{(X,T,Y)\sim P_{X}\times\tilde{P}_{T|X}\times P_{Y|T,X}}\{Y\in\hat{C}(X,T)\}\leq 1-\alpha+\Delta_{w}+cn^{\frac{1}{r-1}}$ (34)
Proof.
We can reformulate our problem as a covariate shift scenario by treating the treatment variable as part of the covariates, i.e., defining $X^{*}=[X,T]$. Under this transformation:
  • The proof for setting S.1 follows directly from Theorem 2 in tibshirani_conformal_2019.
  • The proof for setting S.2 aligns with Proposition 1 in lei_conformal_2021. While their work focuses explicitly on split-weighted conformalized quantile regression (CQR) (romano_conformalized_2019), the argument extends to WCP because it only depends on the weighted exchangeability of nonconformity scores and the boundedness of the likelihood ratio function.
  • Similarly, the proof for setting S.3 follows from Theorem 3 in lei_conformal_2021, along with its corresponding derivation.
Remark 1.
$r$ specifies which moment of the likelihood ratio $w(X,T)$ is being considered. Larger $r$ corresponds to stricter regularity conditions on $w(X,T)$. $M_{r}$ defines the upper bound on the $r$-th moment of $w(X,T)$, ensuring the likelihood ratio does not grow too large and remains well-behaved.
Remark 2.
Note that the term $cn^{\frac{1}{r-1}}$, represents the upper bound of the expectation of maximum weight (probability), i.e., $\mathbb{E}\left[\max_{i\in[1,...,n]\cup\{\infty\}}p^{w}_{i}(X_{n+1})\right]$, which under no covariate shift is equal to $\frac{1}{n+1}$ the upper bound of unweighted conformal prediction.
Remark 3.
The bounding condition assumed in S.2 and S.3 in Proposition 1, $(\mathbb{E}[w(X,T)^{r}])^{\frac{1}{r}}\leq M_{r}<\infty$, that $\mathbb{E}[w(X,T)^{r}]<\infty$ implies that $\mathbb{P}_{(X,T)\sim P_{X}\times P_{T|X}}(w(X)<\infty)=1$ and $\mathbb{E}[w(X)]<\infty$ (lei_conformal_2021), i.e. $P_{X}\times\tilde{P}_{T|X}$ is absolutely continuous with respect to $P_{X}\times P_{T|X}$.
Remark 4.
For setting S.1, the overlap or positivity assumption can be violated, i.e., $\frac{d\tilde{P}_{T|X}}{dP_{T|X}}=\infty$ in terms of the interventional distribution. However, this results in the trivial interval $(-\infty,\infty)$, since $w(X_{i})=0,\forall i\in[1,...,n]$ and $w(X_{n+1})=\infty$ resulting in $p^{w}_{i}(X_{n+1})=0,\forall i\in[1,...,n]$ and $p^{w}_{n+1}=1$.
Remark 5.
Since inductive (or split) conformal prediction is a special case of conformal prediction, Proposition 1 also applies to inductive conformal prediction, which we use in our experiments.
Remark 6.
With an estimated likelihood ratio under weighted CQR, our approach also follows the asymptotic double robustness result (see Theorem 1 (lei_conformal_2021)).
Appendix B Algorithm pseudocode and computational analysis
B.1 Propensity-based Weighted Conformal Prediction Pseudocode
Algorithm 1 presents the fit procedure for both the Local and the Global Propensity WCP, using their respective weights $w_{l,p}^{t}$ and $w_{g,p}^{t}$ for an array of treatment values we want to evaluate $t_{eval}$, assuming a uniform interventional distribution. The pseudocode is written for any Kernel, although in the experiments, we used the Gaussian kernel as presented in the methodology section. The pseudocode assumes either a pre-fitted propensity estimator $\hat{\pi}$ or having access to an Oracle estimator. The method used to fit the propensity estimator in this paper is presented in Appendix B.2. Algorithm 2 then presents how the prediction intervals for a significance level $\alpha$ are generated using both Local and Global Propensity WCP as the implementation is the same for both methods. The get_interval function is the prediction interval function of the WCP method.
Algorithm 1 Fit and calibrate Local or Global Propensity WCP
1:Input: Training covariates $X_{tr}$, calibration covariates $X_{cal}$, training outcome $y_{tr}$, calibration outcome $y_{cal}$, training treatment values $T_{tr}$, calibration treatment values $T_{cal}$, calibrated PropensityEstimator or oracle $\hat{\pi}$, to evaluate treatments in array $t_{eval}$, kernel $K$, CADRF learner $\hat{\mu}$
2: Fit CADRF $\hat{\mu}$ on $(X_{tr},T_{tr})$ to predict $y_{tr}$
3: Calculate propensities $\pi_{cal}=\hat{\pi}(X_{cal})$
4:if Global Propensity WCP then
5:  Calculate weights: $w_{g,p}=1/\pi_{cal}$
6:  Define WCPWCP as Weighted Conformal Prediction with learner $\hat{\mu}$ and weights $w_{g,p}$ on $(X_{cal},T_{cal},y_{cal})$
7:  Calibrate WCPWCP
8:else if Local Propensity WCP then
9:  for $t$ in $t_{eval}$ do
10:   Calculate weights: $w_{l,p}^{t}=K(T_{cal},t)/\pi_{cal}$
11:   Define $WCP_{t}$ as Weighted Conformal Prediction with learner $\hat{\mu}$ and weights $w_{l,p}^{t}$ on $(X_{cal},T_{cal},y_{cal})$
12:   Calibrate $WCP_{t}$
13:  end for
14:end if
15:Output: Calibrated models $\{WCP_{t}:t\in t_{\text{eval}}\}$ for Local Propensity WCP or WCPWCP for Global Propensity WCP
Algorithm 2 Provide uncertainty estimates Local and Global Propensity WCP
1:Input: Test sample $X_{n+1}$, calibrated PropensityEstimator or oracle $\hat{\pi}$, $k$ to evaluate treatments in array $t_{eval}$, kernel $K$, CADRF learner $\hat{\mu}$, calibrated $WCP_{t}$ for all $t$ in $t_{eval}$, significance $\alpha$
2: Calculate $\pi_{n+1}=\hat{\pi}(X_{n+1})$
3: Calculate weights $w=1/\pi_{cal}$
4:for $t$ in $t_{eval}$ do
5:  Predict outcome: $\hat{\mu}(X_{n+1},t)$
6:  Obtain prediction interval: $\hat{C}_{n+1}^{t}=\text{get\_interval}(WCP_{t},(X_{n+1},t),\alpha,w^{t})$
7:end for
8:Output: Prediction intervals $\big[\hat{C}_{n+1,\alpha}^{t_{\text{eval},1}},\dots,\hat{C}_{n+1,\alpha}^{t_{\text{eval},k}}\big]$
B.2 Propensity Distribution Estimation Pseudocode
@@
Algorithm 3 presents the propensity distribution estimation using Conformal Predictive Systems (CPS). This results in a propensity distribution array $\pi_{arr}$ with the calculated propensity density for each sample in $X_{cal}$. expexp is the exponential function and $len(X)$ denotes the length of the array $X$.
@@
1:Input: Training covariates $X_{tr}$, calibration covariates $X_{cal}$, training treatment values $T_{tr}$, calibration treatment values $T_{cal}$, Kernel Density Estimator KDKD
@@
3: Calibrate CPS using $X_{cal}$ and $T_{cal}$
4: Initialize $\pi_{arr}$ as an array of length $len(X_{cal})$
@@
9:Output: Propensity array $\pi_{\text{arr}}$
B.3 Computational Overhead
The computational overhead is greatest for Local Propensity WCP due to the evaluation over multiple treatment values, so we will focus on this version. Let $m$ denote the number of treatment values in the evaluation array $t_{\text{eval}}$. In this case, the computational overhead compared to standard weighted conformal prediction (WCP) scales linearly with the number of treatment values, i.e., $O(m\cdot\text{WCP})$, where WCP refers to the cost of standard weighted conformal prediction. In addition, calculating the propensities $\pi_{\text{cal}}$ on the calibration set incurs an additional computational cost, which depends on the size of the calibration set and the chosen propensity estimator. This step can be done once beforehand, so it does not need to be repeated during each evaluation.
If the treatment values in $t_{\text{eval}}$ are known and fixed, the calibration for each treatment value can be precomputed and stored, resulting in saved $WCP_{t}$ models. This means that, during inference, the computational overhead is reduced to calculating the propensity for a single new sample once and performing $m$ predictions using the CADRF, followed by retrieving the prediction intervals for each treatment value using the pre-calibrated $WCP_{t}$. Thus, the inference overhead is $O(m)$ for a single inference, consisting of a propensity calculation and $m$ predictions and interval retrievals. In the case of a non-static or on-demand $t_{\text{eval}}$, the overhead is additive as we need $O(m\dot{W}CP)$ calibrations and directly afterward $O(m)$ for the inference.
If there is no Oracle propensity estimator, we need to fit the propensity estimator, which, in our case, also involves fitting the Kernel Density Estimator (KDE) for each sample in $X_{\text{cal}}$, as detailed in Algorithm 3. This introduces an extra layer of computational overhead, which depends on the size of the calibration set and the output of the CPS, which is an empirical distribution of the treatment values for $x_{cal}$. The KDE fitting step needs to be performed for each element of $X_{\text{cal}}$, resulting in a complexity of $O(\text{len}(X_{\text{cal}})\cdot\text{KDE})$, where $\text{len}(X_{\text{cal}})$ is the number of calibration samples and KDE denotes the cost of fitting the KDE.
@@ B.3 Computational Overhead
In comparison to the work of schroder_conformal_2024, our approach differs in several key aspects. First, the aim of their work is different from ours. The aim of schroder_conformal_2024 is to provide prediction intervals for the causal effect of treatment interventions where the treatment value is continuous. In our work, the goal is to provide prediction intervals for dose-response models instead of treatment interventions, answering a different causal question. However, adjusting our work to interventions is possible; In the case of soft interventions, the target distribution propensity changes and thus substituting the current uniform distribution in the weights $w(x)$ with the new target propensity distribution covers the soft intervention case. For hard interventions, this is an evaluation for a single treatment value which is similar to the local propensity method, but for only that target treatment value. Secondly, their approach differs in their conformal prediction approach where they want to provide correct prediction intervals for a single sample, single $\alpha$ value, and single treatment using a mathematical solver based on the proposed weighted conformal prediction by gibbs_adaptive_2021. Thirdly, they frame the propensity or covariate shift differently as either a Dirac distribution for a hard intervention, or a different propensity distribution in the case of a soft intervention. This is a direct consequence of their aim to quantify the causal effect of a single intervention, compared to providing a dose-response model in our case which requires a uniform assumption. Fourthly, the experimental setup of schroder_conformal_2024 does not address the impact of a treatment covariate shift as shown by Figure 6 and Figure 7, where even standard conformal prediction (CP) achieves the required empirical coverage. Lastly, we also approach the propensity estimation in cases with unknown propensity as an uncertainty quantification problem and tackle it with conformal predictive systems. In the end, our approach offers a different solution on continuous treatment effects through dose-response modelling.
@@ Appendix D Additional Results
(a) Setup = 2, Scenario = 1
(b) Setup = 2, Scenario = 2
(c) Setup = 3, Scenario = 1
(d) Setup = 3, Scenario = 2
Figure 16: Plot of the CADRF RMSE with $\pm$ RMSE standard deviation across all repeated experiments for the considered treatment values for setup 2 and setup 3. As All WCP and CP methods use the same fitted base CatBoost CADRF learner they are represented by ”CP and WCP”.
(a) Setup = 1, Scenario = 1
(b) Setup = 1, Scenario = 2
(c) Setup = 1, Scenario = 3
(d) Setup = 1, Scenario = 4
Figure 17: Plot of the CADRF RMSE with $\pm$ RMSE standard deviation across all repeated experiments for the considered treatment values for setup 1, scenarios 1 to 4. As All WCP and CP methods use the same fitted base CatBoost CADRF learner they are represented by ”CP and WCP”.
(a) Setup = 1, Scenario = 5
(b) Setup = 1, Scenario = 6
(c) Setup = 1, Scenario = 7
(d) Setup = 1, Scenario = 8
Figure 18: Plot of the CADRF RMSE with $\pm$ RMSE standard deviation across all repeated experiments for the considered treatment values for setup 1, scenarios 5 to 8. As All WCP and CP methods use the same fitted base CatBoost CADRF learner they are represented by ”CP and WCP”.
Figure 19: Bar plot of the coverage across all 50 experiments of the benchmarked methods for setup 1, scenario 1.
Figure 20: Bar plot of the coverage across all 50 experiments of the benchmarked methods for setup 1, scenario 2.
Figure 21: Bar plot of the coverage across all 50 experiments of the benchmarked methods for setup 1, scenario 3.
Figure 22: Bar plot of the coverage across all 50 experiments of the benchmarked methods for setup 1, scenario 4.
Figure 23: Bar plot of the coverage across all 50 experiments of the benchmarked methods for setup 1, scenario 5.
Figure 24: Bar plot of the coverage across all 50 experiments of the benchmarked methods for setup 1, scenario 6.
Figure 25: Bar plot of the coverage across all 50 experiments of the benchmarked methods for setup 1, scenario 7.
Figure 26: Bar plot of the coverage across all 50 experiments of the benchmarked methods for setup 1, scenario 8.
Figure 27: Bar plot of the coverage across all 50 experiments of the benchmarked methods for setup 2, scenario 1.
Figure 28: Bar plot of the coverage across all 50 experiments of the benchmarked methods for setup 2, scenario 2.
Figure 29: Bar plot of the coverage across all 50 experiments of the benchmarked methods for setup 3, scenario 1.
Figure 30: Bar plot of the coverage across all 50 experiments of the benchmarked methods for setup 3, scenario 2.